# AI as Coordination-Compressing Capital:

### v1

@@ Abstract
Task-based models of AI and labor hold organizational structure fixed, analyzing how technology shifts task assignments within a given firm architecture. Yet emerging evidence shows firms flattening hierarchies in response to AI adoption—a phenomenon these models cannot generate. We extend the task-based framework by introducing agent capital ($K_{A}$): AI systems that reduce coordination costs within organizations, expanding managerial spans of control and enabling endogenous task creation. We derive five propositions characterizing how coordination compression affects output, hierarchy, manager demand, wage dispersion, and the task frontier. The model generates a regime fork: depending on whether agent capital complements all workers broadly (general infrastructure) or high-skill managers disproportionately (elite complementarity), the same technology produces either broad-based productivity gains or superstar concentration, with sharply divergent distributional consequences. Numerical simulations with heterogeneous managers and workers across a $2\times 2$ parameter space (elite complementarity $\times$ endogenous task creation) confirm sharp regime divergence: in settings where coordination compression substantially expands employment, economy-wide inequality falls in all regimes, but the rate of reduction is regime-dependent and the manager–worker wage gap widens universally. The distributional impact of AI hinges not on the technology itself but on the elasticity of organizational structure—and on who controls that elasticity.
@@ 1 Introduction
The dominant framework for analyzing AI’s labor market effects is task-based: technology changes which tasks are feasible, who performs them, and at what cost. Acemoglu and Restrepo (2019) formalize this as a displacement-reinstatement dynamic, where automation removes workers from existing tasks while new task creation reinstates labor demand. Agrawal et al. (2026) refine this approach, showing that AI can enhance worker productivity without automating tasks—by expanding the technologically feasible task set ($\bar{\tau}$) or broadening individual workers’ feasible task sets ($\Omega(i)$). Their key insight is that productivity (captured by geometric-mean uniformity of the task distribution) and inequality (captured by Shannon entropy) are different functionals of the same equilibrium distribution and can move independently or even in opposite directions.
@@
Two recent papers embed AI in Garicano’s (2000) knowledge hierarchy but hold organizational costs fixed. Ide and Talamás (2025) show that autonomous AI benefits the most knowledgeable while non-autonomous AI benefits the least knowledgeable, but their communication cost parameter is exogenous—AI reshuffles occupational roles within an unchanged organizational architecture. Xu et al. (2025) show that generative AI can paradoxically narrow spans of control through induced deskilling, but again hold communication costs fixed while varying the knowledge partition between workers and managers. Cheng et al. (2025) model AI replacement in teams with sequential information flow, showing that positional structure—not just skill level—determines displacement risk. Our approach is complementary: we hold the knowledge partition fixed and vary the coordination cost itself. This is precisely the channel that Garicano and Rossi-Hansberg (2006) showed has first-order effects on inequality—reductions in communication costs widen spans, compress hierarchies, and concentrate returns at the top—but that neither paper endogenizes.
This paper makes three contributions. First, we introduce agent capital ($K_{A}$) as a distinct production input that reduces coordination costs within organizations. Unlike task-level AI, which changes what workers can do, agent capital changes how work is organized—compressing the coordination friction $c(K_{A})$ that determines spans of control, hierarchy depth, and the boundary between managerial and production labor. Second, we derive five formal propositions characterizing the comparative statics of coordination compression: output rises (Proposition 1), spans of control expand (Proposition 2), manager demand falls (Proposition 3), wage dispersion rises under elite complementarity (Proposition 4), and the task frontier expands endogenously (Proposition 5). Third, these propositions generate a regime fork. When agent capital is broadly accessible (low $\beta$), coordination compression produces wage compression and broad-based productivity gains—a “general infrastructure” equilibrium. When agent capital disproportionately amplifies high-skill managers (high $\beta$), the same technology produces superstar concentration—an “elite complementarity” equilibrium. The distributional consequences of AI depend critically on which regime obtains.
@@ 2.2 Span of Control
as the maximum team size manager $i$ can effectively coordinate, given the technology of coordination. This is a technological parameter—a capacity constraint—not the solution to a manager’s optimization problem. Note that $L_{\text{eff},i}=Q_{i}/(1+c_{i}n_{i})$ embodies a tension: adding a worker of skill $q_{j}$ raises the numerator $Q_{i}$ by $q_{j}$ but also raises the coordination penalty $c_{i}n_{i}$ in the denominator. A worker contributes positively on net only when $q_{j}$ is large enough relative to the existing coordination burden. [^1] An equivalent micro-foundation: each coordination link consumes $c_{i}$ units of the manager’s unit time endowment, yielding the constraint $c_{i}\cdot n_{i}\leq 1$, i.e., $n_{i}\leq 1/c_{i}=S_{i}$. This capacity constraint, rather than diminishing returns in $L_{\text{eff}}$, is what bounds team size. In the homogeneous case, $S=(1+\gamma K_{A})/c_{0}$, which is linear in $K_{A}$ and increasing in $\gamma$. In the newsroom: at baseline ($K_{A}=0$), a section editor with $c_{0}=0.3$ can coordinate $S=1/0.3\approx 3.3$ reporters. At $K_{A}=5$ with $\gamma=1$, capacity rises to $S=(1+5)/0.3=20$. The editor-in-chief can now oversee the entire reporting staff directly—section editors become redundant.
@@ 3 Propositions
In a multi-layer hierarchy with homogeneous span $S>1$, total managers are $M=N/S+N/S^{2}+\cdots=N/(S-1)$. Since $S=(1+\gamma K_{A})/c_{0}$ and $c_{0}<1$ ensures $S>1$ for all $K_{A}\geq 0$, we have $\partial M/\partial K_{A}=-N\gamma/[c_{0}(S-1)^{2}]<0$. In the heterogeneous case, total supervisory capacity $D=\sum S_{i}$ is increasing in $K_{A}$ (since each $\partial S_{i}/\partial K_{A}>0$ by Proposition 2): each manager can coordinate more workers, so fewer managers are needed for a given workforce. Under $\beta>1$, the reduction is compositionally concentrated: top managers absorb disproportionately more workers while bottom managers become redundant. ∎
@@
Under proportional worker allocation, the wage ratio between managers $i$ and $j$ reduces to $R=[(1+\gamma K_{A}s_{i}^{\beta})/(1+\gamma K_{A}s_{j}^{\beta})]^{\alpha}$. For $s_{i}>s_{j}$ and $\beta>0$: $dR/dK_{A}=\alpha\cdot u^{\alpha-1}\cdot\gamma(s_{i}^{\beta}-s_{j}^{\beta})/(1+\gamma K_{A}s_{j}^{\beta})^{2}>0$, where $u=(1+\gamma K_{A}s_{i}^{\beta})/(1+\gamma K_{A}s_{j}^{\beta})$, so all pairwise wage ratios are strictly increasing in $K_{A}$. At $K_{A}=0$, all wages are equal (Gini =0=0). For $K_{A}>0$, normalize wages by the bottom manager’s wage: $\hat{w}_{i}=(S_{i}/S_{1})^{\alpha}$. Since $\hat{w}_{1}=1$ is fixed while every $\hat{w}_{i}$ for $i>1$ is strictly increasing in $K_{A}$, the cumulative share of the bottom $k$ managers, $\sum_{i=1}^{k}\hat{w}_{i}/\sum_{i=1}^{n}\hat{w}_{i}$, is strictly decreasing for all $k<n$: the denominator grows strictly faster than any partial numerator sum. Therefore the Lorenz curve at $K_{A}^{\prime}>K_{A}>0$ lies strictly below the Lorenz curve at $K_{A}$, and the Gini rises monotonically. For $\beta>1$, the mapping $s\mapsto s^{\beta}$ is convex, amplifying skill differences superlinearly (“Winner Takes All”); for $\beta\in(0,1)$, the mapping is concave and the Gini increase is modest (“Gentle Compression”). ∎
@@
Part 1: $dT/dK_{A}=T_{0}\delta>0$, so the frontier is strictly increasing. Part 2: Employment $E=\min(D,T,N)$ where $D=\sum S_{i}$ is total supervisory capacity. At low $K_{A}$, the economy is demand-constrained ($E=D$, increasing by Proposition 2). As $K_{A}$ rises and $\delta>0$, the frontier ceiling also rises, so $E$ increases whether the binding constraint is capacity or the frontier. Since $D$, $T$, and $N$ are continuous in $K_{A}$, $E=\min(D,T,N)$ is continuous, ensuring no downward jumps at transitions between constraint regimes. The unemployment rate u=1E/Nu=1-E/N is weakly decreasing. The parameter $\delta$ determines whether coordination compression creates fixed-pie reallocation ($\delta=0$: employment hits the ceiling $T_{0}$) or expanding-pie absorption ($\delta>0$: new tasks absorb displaced labor). ∎
@@ 4.1 The $2\times 2$ Taxonomy
Agent capital amplifies top managers disproportionately. No new tasks absorb displaced workers. Superstar managers scale across larger spans while others stagnate. Output rises—but inequality rises sharply. This extends Rosen’s (1981, 1982) superstar mechanism with a coordination channel.
@@ 4.4 Economy-Wide Distributional Results
Second, the manager–worker wage gap rises in all regimes. For this calibration, the mean manager earns approximately $1.8\times$ the mean employed worker at $K_{A}=0$, rising to approximately $5.4\times$ at $K_{A}=10$ under Gentle Compression. The direction is not regime-dependent: coordination compression universally concentrates returns in the coordinating layer, because managers capture a fixed share of team output ($(1-\alpha)Y_{i}$) while the worker share ($\alpha Y_{i}$) is divided among all team members. As spans expand, each manager’s output grows while workers share among more colleagues.
@@
Results are qualitatively robust to random (non-assortative) worker assignment; PAM is the primary specification, motivated by Garicano and Rossi-Hansberg (2006).
@@ 5.1 Testable Predictions
The model generates testable organizational predictions. Propositions 23 predict declining management share and expanding spans of control as coordination-compressing AI tools diffuse—patterns consistent with Ewens and Giroud’s (2025) finding that firms flattened hierarchies after AI adoption. The critical empirical question is how AI coordination adoption covaries with managerial skill, which determines $\beta$ and thus the regime. These constructs are operationalizable via O*NET task taxonomies and AI exposure indices, though causal identification of coordination compression effects remains an open design challenge.
@@ 5.3 $\beta$ as the Policy Lever
A natural question is whether $\beta$ could be negative—whether AI might help low-skill coordinators disproportionately, leveling rather than amplifying skill differences. Empirical evidence supports this for production tasks: Brynjolfsson et al. (2025) find that generative AI narrows the performance gap between novice and expert customer service agents. But coordination tasks differ from production tasks in a critical respect. Production tasks are often codable—AI can provide expert-level scripts to novices. Coordination tasks require judgment under ambiguity: deciding what to delegate, monitoring heterogeneous outputs, resolving conflicts between workstreams. AI provides the capacity to handle more coordination links, but the quality of coordination decisions remains skill-dependent. The model restricts $\beta\geq 0$ because coordination complementarity (AI amplifies coordination skill) is more plausible than coordination substitution (AI replaces coordination judgment). If $\beta$ were negative, the model would predict active skill-leveling among managers—an even stronger version of the general-infrastructure case that would further compress inequality.

### v3

@@ Abstract
Task-based models of AI and labor hold organizational structure fixed. We introduce agent capital: AI that reduces coordination costs, expanding spans of control and enabling endogenous task creation. Five propositions characterize how coordination compression affects output, hierarchy, manager demand, wage dispersion, and the task frontier. The model generates a regime fork: the same technology produces broad-based gains or superstar concentration depending on who benefits from coordination compression. Simulations with heterogeneous workers confirm sharp regime divergence. Within the simulated firm, inequality falls in all regimes as coordination compression expands employment, but the manager–worker wage gap widens universally. The distributional impact hinges on who controls organizational elasticity.
@@ 1 Introduction
The dominant framework for analyzing AI’s labor market effects is task-based: technology changes which tasks are feasible, who performs them, and at what cost. Acemoglu and Restrepo (2019) formalize this as a displacement-reinstatement dynamic, where automation removes workers from existing tasks while new task creation reinstates labor demand. Agrawal et al. (2026) refine this approach, showing that AI can enhance worker productivity without automating tasks—by expanding the technologically feasible task set ($\bar{\tau}$) or broadening individual workers’ feasible task sets ($\Omega(i)$). Their key insight is that productivity (captured by geometric-mean uniformity of the task distribution) and inequality (captured by Shannon entropy, which measures distributional uniformity and thus moves inversely with inequality) are different functionals of the same equilibrium distribution and can move independently or even in opposite directions.
@@
Recent papers embed AI in Garicano’s (2000) knowledge hierarchy but hold organizational costs fixed. Ide and Talamás (2025) show that autonomous AI benefits the most knowledgeable while non-autonomous AI benefits the least knowledgeable, but their communication cost parameter is exogenous—AI reshuffles occupational roles within an unchanged organizational architecture. Xu et al. (2025) show that generative AI can paradoxically narrow spans of control through induced deskilling, but again hold communication costs fixed while varying the knowledge partition between workers and managers. Cheng et al. (2025) model AI replacement in teams with sequential information flow, showing that positional structure—not just skill level—determines displacement risk. Our approach is complementary: we hold the knowledge partition fixed and vary the coordination cost itself. This is precisely the channel that Garicano and Rossi-Hansberg (2006) showed has first-order effects on inequality—reductions in communication costs widen spans, compress hierarchies, and concentrate returns at the top—but that neither Ide and Talamás (2025) nor Xu et al. (2025) parameterizes as a function of AI capability.
This paper makes three contributions. First, we introduce agent capital ($K_{A}$) as a distinct production input that reduces coordination costs within organizations. Unlike task-level AI, which changes what workers can do, agent capital changes how work is organized—compressing the coordination friction $c(K_{A})$ that determines spans of control, hierarchy depth, and the boundary between managerial and production labor. Second, we derive five formal propositions characterizing the comparative statics of coordination compression: output rises (Proposition 1), spans of control expand (Proposition 2), manager demand falls (Proposition 3), wage dispersion rises under elite complementarity (Proposition 4), and the task frontier expands endogenously (Proposition 5). Third, these propositions generate a regime fork. When agent capital is broadly accessible (low $\beta$), coordination compression produces broad-based productivity gains and, within the firm, reduces inequality through employment expansion—a “general infrastructure” equilibrium. When agent capital disproportionately amplifies high-skill managers (high $\beta$), the same technology produces superstar concentration—an “elite complementarity” equilibrium. The distributional consequences of AI depend critically on which regime obtains.
@@ 2.2 Span of Control
as the maximum team size manager $i$ can effectively coordinate, given the technology of coordination. This is a technological parameter—a capacity constraint—not the solution to a manager’s optimization problem. Two closures are possible: one could treat team size as an interior choice where the manager trades off the marginal worker’s quality contribution against coordination drag in $L_{\text{eff}}$, or one could impose the capacity constraint $n_{i}\leq S_{i}$ as binding. We adopt the latter throughout. Specifically, we assume labor is sufficiently abundant that each manager’s team reaches the capacity constraint $S_{i}$. This holds when the total worker pool $N$ exceeds total supervisory demand $\sum S_{i}$, or equivalently when outside options are low enough that workers prefer team assignment at the prevailing wage share. Under this closure, $L_{\text{eff},i}$ in (3) gives team output conditional on $n_{i}=S_{i}$; spans are set by coordination technology, not by an interior optimality condition. Note an important implication: when $n_{i}=S_{i}=1/c_{i}$, the denominator of (3) becomes $1+c_{i}/c_{i}=2$, so $L_{\text{eff},i}=Q_{i}/2$ regardless of $c_{i}$. At the binding capacity constraint, coordination compression raises output not by reducing friction for a fixed team, but by expanding the team itself: lower $c_{i}$ increases $S_{i}$, admitting more workers, which raises $Q_{i}$. Proposition 1’s “holding team allocations fixed” qualifier is essential—it establishes the output effect for a given team size, while the capacity-constraint closure channels the practical effect through team expansion. [^1] An equivalent micro-foundation: each coordination link consumes $c_{i}$ units of the manager’s unit time endowment, yielding the constraint $c_{i}\cdot n_{i}\leq 1$, i.e., $n_{i}\leq 1/c_{i}=S_{i}$. This capacity constraint, rather than diminishing returns in $L_{\text{eff}}$, is what bounds team size. In the homogeneous case, $S=(1+\gamma K_{A})/c_{0}$, which is linear in $K_{A}$ and increasing in $\gamma$. In the newsroom: at baseline ($K_{A}=0$), a section editor with $c_{0}=0.3$ can coordinate $S=1/0.3\approx 3.3$ reporters. At $K_{A}=5$ with $\gamma=1$, capacity rises to $S=(1+5)/0.3=20$. The editor-in-chief can now oversee the entire reporting staff directly—section editors become redundant.
@@ 3 Propositions
In a multi-layer hierarchy with homogeneous span $S>1$ and $L$ layers, total managers are $M=\sum_{\ell=1}^{L}N/S^{\ell}$. For large $N$ (so that $L\to\infty$), this converges to $M\approx N/(S-1)$; we use this limiting approximation for tractability. Since $S=(1+\gamma K_{A})/c_{0}$ and $c_{0}<1$ ensures $S>1$ for all $K_{A}\geq 0$, we have $\partial M/\partial K_{A}=-N\gamma/[c_{0}(S-1)^{2}]<0$. The multi-layer formula extends beyond the two-layer production model of Section 2; we include it to show that the comparative static generalizes to deeper hierarchies. In the two-layer case that the production model and simulation directly implement, the result is more immediate: total supervisory capacity $D=\sum S_{i}$ is increasing in $K_{A}$ (since each $\partial S_{i}/\partial K_{A}>0$ by Proposition 2), so fewer managers are needed to cover a given workforce. Under $\beta>1$, the reduction is compositionally concentrated: top managers absorb a larger share of workers while bottom managers become redundant. (The simulation holds the manager pool fixed at $n=20$, so Proposition 3 manifests as excess managerial capacity rather than explicit displacement.) ∎
@@
Under proportional worker allocation—each manager receives $n_{i}=\lambda S_{i}$ workers (for common $\lambda$) drawn from a pool with equal average skill, so that team quality satisfies $Q_{i}/Q_{j}=S_{i}/S_{j}$—the wage ratio between managers $i$ and $j$ reduces to $R=[(1+\gamma K_{A}s_{i}^{\beta})/(1+\gamma K_{A}s_{j}^{\beta})]^{\alpha}$. (Under PAM, where high-skill workers are matched to high-span managers, $Q_{i}$ may grow faster than $S_{i}$ for top managers, reinforcing the dispersion result.) For $s_{i}>s_{j}$ and $\beta>0$: $dR/dK_{A}=\alpha\cdot u^{\alpha-1}\cdot\gamma(s_{i}^{\beta}-s_{j}^{\beta})/(1+\gamma K_{A}s_{j}^{\beta})^{2}>0$, where $u=(1+\gamma K_{A}s_{i}^{\beta})/(1+\gamma K_{A}s_{j}^{\beta})$, so all pairwise wage ratios are strictly increasing in $K_{A}$. At $K_{A}=0$, all wages are equal (Gini =0=0). For $K_{A}>0$, normalize wages by the bottom manager’s wage: $\hat{w}_{i}=(S_{i}/S_{1})^{\alpha}$. Since $\hat{w}_{1}=1$ is fixed while every $\hat{w}_{i}$ for $i>1$ is strictly increasing in $K_{A}$, the Lorenz share of the bottom $k$ managers is $L_{k}=\sum_{i=1}^{k}\hat{w}_{i}/\sum_{i=1}^{n}\hat{w}_{i}$. For $k<n$, we show $dL_{k}/dK_{A}<0$. By the quotient rule, $dL_{k}/dK_{A}<0$ iff $(\sum_{i=1}^{k}\hat{w}_{i}^{\prime})(\sum_{i=1}^{n}\hat{w}_{i})<(\sum_{i=1}^{n}\hat{w}_{i}^{\prime})(\sum_{i=1}^{k}\hat{w}_{i})$, where primes denote $K_{A}$-derivatives. The pairwise ratio result (above) implies that $\hat{w}_{i}^{\prime}/\hat{w}_{i}$ is strictly increasing in $i$: managers with larger $s_{i}^{\beta}$ have strictly higher proportional wage growth rates. Since the bottom $k$ managers have the lowest growth rates, the weighted average $\sum_{i=1}^{k}\hat{w}_{i}^{\prime}/\sum_{i=1}^{k}\hat{w}_{i}$ is strictly less than $\sum_{i=1}^{n}\hat{w}_{i}^{\prime}/\sum_{i=1}^{n}\hat{w}_{i}$, which gives the required inequality. Therefore the Lorenz curve at $K_{A}^{\prime}>K_{A}>0$ lies strictly below the Lorenz curve at $K_{A}$, and the Gini rises monotonically. For $\beta>1$, the mapping $s\mapsto s^{\beta}$ is convex, amplifying skill differences superlinearly (“Winner Takes All”); for $\beta\in(0,1)$, the mapping is concave and the Gini increase is modest (“Gentle Compression”). ∎
Remark. The wage dispersion result depends on the Cobb-Douglas rent-sharing convention ($w_{M}=(1-\alpha)Y_{i}$, worker shares proportional to quality). Under alternative wage determination (e.g., Nash bargaining over team surplus), the magnitude and potentially the direction of this result may differ. The qualitative finding that differential span expansion amplifies managerial wage heterogeneity is more general than the specific Gini characterization.
@@
Part 1: $dT/dK_{A}=T_{0}\delta>0$, so the frontier is strictly increasing. Part 2: Employment $E=\min(D,T,N)$ where $D=\sum S_{i}$ is total supervisory capacity. At low $K_{A}$, the economy is demand-constrained ($E=D$, increasing by Proposition 2). As $K_{A}$ rises and $\delta>0$, the frontier ceiling also rises, so $E$ increases whether the binding constraint is capacity or the frontier. Since $D$, $T$, and $N$ are continuous in $K_{A}$, $E=\min(D,T,N)$ is continuous, ensuring no downward jumps at transitions between constraint regimes. The unemployment rate u=1E/Nu=1-E/N is weakly decreasing. The parameter $\delta$ determines whether coordination compression creates fixed-pie reallocation ($\delta=0$: employment hits the ceiling $T_{0}$) or expanding-pie absorption ($\delta>0$: new tasks absorb displaced labor). We note that this employment result is, by construction, a consequence of the monotonicity assumptions on $D$ and $T$; it characterizes the model’s mechanism rather than delivering a surprising prediction. The economy-wide inequality reduction reported in Section 4.4 follows from this extensive-margin expansion (moving workers from zero to positive income) and should be interpreted as such. ∎
@@ 4.1 The $2\times 2$ Taxonomy
Agent capital amplifies top managers selectively. No new tasks absorb displaced workers. Superstar managers scale across larger spans while others stagnate. Output rises—but inequality rises sharply. This extends Rosen’s (1981, 1982) superstar mechanism with a coordination channel.
@@ 4.4 Economy-Wide Distributional Results
Second, the manager–worker wage gap rises in all regimes. For this calibration, the mean manager earns approximately $1.8\times$ the mean employed worker at $K_{A}=0$, rising to approximately $5.4\times$ at $K_{A}=10$ under Gentle Compression. The direction is not regime-dependent: coordination compression universally concentrates returns in the coordinating layer, because managers capture a fixed share of team output ($(1-\alpha)Y_{i}$) while the worker share ($\alpha Y_{i}$) is divided among all team members. As spans expand, each manager’s output grows while workers share among more colleagues. This result is a direct consequence of the Cobb–Douglas rent-sharing convention: the manager captures $(1-\alpha)$ of a growing pie while the worker share is diluted across an expanding team. Under alternative wage determination (e.g., Nash bargaining or competitive labor markets with outside options), the gap trajectory could differ. The qualitative mechanism—that coordination compression amplifies the organizational leverage of the coordinating layer—is more general than the specific magnitude.
@@
Results are qualitatively robust to random (non-assortative) worker assignment: all five qualitative checks (output rising, manager Gini ordering, economy-wide Gini falling, gap rising, employment ordering) pass under random matching for all four regimes. PAM is the primary specification because it arises as an equilibrium in knowledge hierarchies where higher-skill managers generate larger team surplus and can bid more for higher-skill workers (Garicano and Rossi-Hansberg, 2006). In the present model, PAM is imposed rather than derived from a competitive assignment game, and the regime fork’s qualitative structure does not depend on it. PAM sharpens the magnitudes—particularly the Gini levels under high $\beta$—but the directional results are driven by coordination compression, not the assignment rule.
@@ 5.1 Testable Predictions
The model generates testable organizational predictions. Propositions 23 predict declining management share and expanding spans of control as coordination-compressing AI tools diffuse—patterns documented by Ewens and Giroud’s (2025) finding that firms flattened hierarchies after AI adoption. The critical empirical question is how AI coordination adoption covaries with managerial skill, which determines $\beta$ and thus the regime. These constructs are operationalizable via O*NET task taxonomies and AI exposure indices, though causal identification of coordination compression effects remains an open design challenge.
@@ 5.3 $\beta$ as the Policy Lever
A natural question is whether $\beta$ could be negative—whether AI might help low-skill coordinators disproportionately, leveling rather than amplifying skill differences. Empirical evidence supports this for production tasks: Brynjolfsson et al. (2025) find that generative AI narrows the performance gap between novice and expert customer service agents. But coordination tasks differ from production tasks in a critical respect. Production tasks are often codable—AI can provide expert-level scripts to novices. Coordination tasks require judgment under ambiguity: deciding what to delegate, monitoring heterogeneous outputs, resolving conflicts between workstreams. AI provides the capacity to handle more coordination links, but the quality of coordination decisions remains skill-dependent. The model restricts $\beta\geq 0$ because coordination complementarity (AI amplifies coordination skill) is more plausible than coordination substitution (AI replaces coordination judgment). If $\beta$ were negative, the mapping $s_{i}\mapsto s_{i}^{\beta}$ would invert the ranking: managers with lower raw skill would receive more effective agent capital ($s_{i}^{\beta}>1$ for $s_{i}<1$ when $\beta<0$). For managers with skills bounded away from zero, this would compress the span distribution relative to $\beta=0$, producing a leveling effect. However, for $s_{i}$ near zero the power function $s_{i}^{\beta}$ diverges, so the distributional implications of $\beta<0$ depend on the lower bound of the skill distribution—a complication the model avoids by restricting $\beta\geq 0$.